OpenStudy (anonymous):

tan(pi/4)-sin(pi/6) How would I find the exact value?

OpenStudy (anonymous):

Can you go through the steps with me?

OpenStudy (amorfide):

you would literally just use a calculator

OpenStudy (amorfide):

and leave it as a fraction or root

rishavraj (rishavraj):

u know the value of tan (pi/4) and sin(pi/6)???

OpenStudy (anonymous):

I can't use a calculator for this part, and tan is 45 degrees and sin is 30 degrees?

OpenStudy (anonymous):

@rishavraj and @amorfide

OpenStudy (amorfide):

then you gotta remember your triangles 'I am useless here I will let rishavraj explain

rishavraj (rishavraj):

see tan 45 = 1 sin 30 = 1/2...u know tht??

OpenStudy (anonymous):

Where are you getting those numbers from?

OpenStudy (anonymous):

Ohh I see

rishavraj (rishavraj):

sin 45 = cos 45 = \[1/\sqrt{2}\] and tan 45 = sin45/cos 45

OpenStudy (anonymous):

I'm sorry I'm not understanding For my final answer I'm supposed to get 1/2

rishavraj (rishavraj):

tan 45 = 1 sin 30 = \[\frac{ 1 }{ 2 }\] plug the values

OpenStudy (anonymous):

I see where the sin30=1/2 comes from but where does the tan45=1 come from?

rishavraj (rishavraj):

http://www.mathsisfun.com/sine-cosine-tangent.html